# Near-real-time flood inundation monitoring by Bayesian analysis for change point problems for Sentinel-1 time series

### v2

@@ 3.2 Bayesian analysis for change point problems
[…]
@@
We implemented BCP using the bcp R package [Erdman2008-kv]. We selected the default tuning parameters of $\gamma=\lambda=0.2$ as recommended by Barry and Hartigan [Barry1993-qp, Erdman2008-kv]. Posterior means and probabilities were calculated based on 500 MCMC iterations after discarding the first 50 as burn-in, a configuration that achieved convergence while maintaining computational efficiency.
@@
3.3 De-noising
Sentinel-1 SAR images inherently contain multiplicative speckle noise which may lead to false detections of flooded areas. To reduce this noise while preserving spatial details, we applied a local median filter to the results of the BCP analysis. A local median filter with a window size of $w$ x $w$ pixels was applied to the BCP output. The window size $w$ is a parameter that can be adjusted based on the spatial resolution of the input imagery and the desired level of smoothing. After applying the filter, the smoothed values were thresholded using a parameter $t$. Pixels with values above this threshold were considered as potential flood areas, while those below were masked out.
To determine optimal parameters for our flood detection method, we conducted a systematic parameter sensitivity analysis focusing on the $w$ and $t$. For the window size parameter, we explored values ranging from 3 to 15 pixels with a step size of 2. The lower bound of 3 represents the minimum window size needed for effective noise reduction while the upper bound of 15 was chosen to avoid excessive smoothing that could eliminate small but significant flooded areas. The 20 m spatial resolution of our analysis data meant that a window size of 3 would cover an area of 60m × 60m on the ground, while a window size of 15 would cover 300 m × 300 m.
For the threshold parameter ($t$), we investigated values between 0.1 and 0.9 with a step size of 0.1. This comprehensive range covers the entire probability space from very permissive detection (0.1) to highly selective identification (0.9). Theoretically, changepoint probabilities above 0.5 indicate a more likely than not change, but in practice, lower thresholds may be effective for detecting subtle hydrological changes. Thus, the parameter optimization process involved evaluating all 63 combinations of window sizes and thresholds (7 window sizes × 9 threshold values) for each flood event (local) or across multiple flood events (global) in diverse geographical settings to ensure the robustness and transferability of our approach.
@@ 4 Evaluation
4.0.1 OTSU Thresholding Method
As a conventional baseline, we implemented the Otsu thresholding method [Otsu1979-dj]. This non-parametric technique determines an optimal threshold by maximizing the between-class variance between flood and non-flood pixels in SAR imagery. We applied the Otsu method to two different input configurations: VV polarization only and VH polarization only.
The Otsu algorithm operates by computing the weighted sum of variances of the two classes:
@@ 4.0.1 OTSU Thresholding Method
where $\omega_{0}$ and $\omega_{1}$ represent the class probabilities, $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ denote the class variances, and $t$ is the threshold value. We followed standard practices in remote sensing literature for the Otsu method, including histogram equalization preprocessing to enhance contrast in the SAR imagery before applying the threshold calculation.
@@ 4.0.2 Siamese CNN Method
The model was trained using binary cross-entropy loss function, with consideration for class imbalance where the positive class weight was calculated to be approximately 35.67. We used the Adam optimizer with an initial learning rate of 0.001 and gradient clipping (clipnorm=1.0), along with a ReduceLROnPlateau scheduler that reduced the learning rate by half after 5 epochs of no improvement. The model was trained for up to 200 epochs with early stopping (patience=20) based on validation loss.
The training dataset comprised 231 image pairs with 58 pairs reserved for validation from urban flooding events. Each input was normalized using Z-score normalization and clipped to ±3 standard deviations to handle outliers. NaN values were replaced with zeros to ensure numerical stability.
@@ 4.0.3 Accuracy assessment
where $N_{TP}$ represents the number of true positives (correctly predicted flood water pixels), $N_{FP}$ represents the number of false positives (incorrectly predicted as flood water), and $N_{FN}$ represents the number of false negatives (flood water pixels not detected).
The F1 score, which provides a balanced measure of the model’s performance, is calculated as:
@@
The Intersection-over-Union (IoU) was aimed to quantify the spatial overlap between the predicted flood water regions and the ground truth annotations:
@@ 5.1 BCP Outputs
Figure 1: Randomly selected patches for: Jubba (Somalia) (a) Bing map that represent non-flood condition, (b) Referenced UrbanSARFloods data set, (c) BCP results used VV input only, (d) BCP results used VH input only, (e) BCP results used VV-VH input; Weihui (China) (f) Bing map that represent non-flood condition, (g) Binary UrbanSARFloods data set, (h) BCP results used VV input only, (i) BCP results used VH input only, (j) BCP results used VV-VH input; NovaKakhovka (Ukraine) (k) Bing map that represent non-flood condition, (l) Binary UrbanSARFloods data set, (m) BCP results used VV input only, (n) BCP results used VH input only, (o) BCP results used VV-VH input.
@@ 5.2 Parameter Sensitivity Analysis
The original BCP results are the posterior probabilities of flood occurrence at each pixel. To delineate the flood extents, we post-processed the BCP results with different parameter settings for threshold and window size and created the binary (flood/non-flood) outputs. Considering the result using VV-VH input only, as this is thought to be the best performance for the BCP, Figure 2 shows the comparison of the binary outputs with different parameter settings for Jubba (Somalia) as an example. The larger $w$ tends to overestimate the flood extent, while the smaller $w$ may fail to omit the speckle noise. The larger $t$ picks up only the flood extent with the higher confidence, while the smaller $t$ may include false positives.
Figure 2: Comparative maps of a randomly selected patch for Jubba (Somalia) with different parameter settings for post-processing.
@@
We identified 44 of 63 parameter combinations that showed no statistically significant difference from the optimal configuration ($p>0.05$). These combinations formed several key clusters: combinations with $t=0.2$ and window sizes 7, 11, and 13 (Mean F1 $>$ 0.59, see table 2); combinations with $t=0.3$ and window sizes 5-15 (Mean F1 $>$ 0.56); and some combinations with $t=0.1$ (Mean F1 $>$ 0.50).
Performance degradation became statistically significant for threshold values $\geq$ 0.4, particularly when combined with extreme window sizes (Figure 3). The decline in performance showed a consistent pattern with increasing threshold values and window size.
Table 2: Top performing parameter combinations and their statistical significance. Selected results only with t=0.2.
@@ 5.3 Comparative Analysis Across Sites
With the use of the optimal parameters for BCP, Figure 4 illustrates the comparison of our method with the conventional and deep-learning approaches across randomly selected three example patches. Each column represents a different study area, with rows showing the referenced UrbanSARFloods data, OTSU thresholding (VV, VH), Siamese CNN, and BCP (VV, VH, VV-VH) with the optimized parameters. The OTSU thresholding approach seems a good performance to Wuhei, but it tends to overestimate the flood extent in Jubba and NovaKakhovka. VV input seems to perform better than VH input in all three sites in terms of noise reduction. Siamese CNN shows a good performance and successfully overcomes the noise compared to the other approaches. However, it tends to underestimate the flood extent in Jubba and NovaKakhovka. BCP with VV or VH alone tends to have noise in the flood extent, while the BCP with VV-VH input shows relatively good in terms of noise reduction.
Figure 4: Flood detection results across randomly selected patches in different geographical contexts and estimated approaches.
@@ 5.4 Quantitative Performance Evaluation
The table LABEL:tab:comparison summarizes the accuracy measures of our approach compared to OTSU thresholding and Siamese CNN approaches. For this evaluation, we used the globally optimazed parameters identified through sensitivity analysis (t=0.2, w=9) with VV-VH input. Our BCP approach achieved F1 scores ranging from 0.41 to 0.76, significantly outperforming OTSU thresholding and Siamese CNN approaches across all sites. The OTSU thresholding approach struggled with false positives, resulting in low precision across all sites. While the Siamese CNN showed improved performance over OTSU, it still fell short of our method’s capabilities.
Table 3: Accuracy assessment of flood detection methods across three flood events.
@@
OTSU (VV) 0.02 0.37 0.03 0.02 0.08 0.49 0.11 0.07 0.03 0.58 0.05 0.03
OTSU (VH) 0.03 0.43 0.04 0.03 0.09 0.57 0.12 0.08 0.03 0.54 0.05 0.03
Siamese-CNN 0.33 0.42 0.34 0.24 0.38 0.26 0.22 0.16 0.18 0.11 0.08 0.05
BCP 0.73 0.79 0.76 0.61 0.57 0.74 0.64 0.48 0.63 0.30 0.41 0.25
@@ 5.5 Analysis of Performance Variations Between Urban and Open Areas
To better understand the performance characteristics across different land cover types, we conducted separate analyses for flood urban and flood open pixels. Tables 4 and 5 present the detailed accuracy metrics for each category and method. In open areas, all methods demonstrated measurable effectiveness in flood detection, though with varying degrees of success. The BCP method achieved the highest performance in Jubba with a precision of 0.37 and recall of 0.75 (IoU: 0.37), followed by strong results in Weihui (precision: 0.31, recall: 0.62, IoU: 0.31). However, its performance declined significantly in NovaKakhovka (precision: 0.05, recall: 0.09, IoU: 0.05). The OTSU method showed consistent but modest performance across sites, with recall values ranging from 0.21 to 0.56, though precision remained low (0.01-0.08). The Siamese CNN demonstrated more balanced performance in open areas, particularly in Weihui where it achieved an IoU of 0.13, though its effectiveness varied considerably across sites.
The analysis of urban environments revealed more challenging conditions for flood detection across all methods. The conventional OTSU method exhibited extremely low precision (0.000-0.002) despite achieving moderate recall values (0.02-0.23), resulting in poor overall performance with F1 scores below 0.003. The Siamese CNN demonstrated notably superior performance in urban areas compared to other methods, achieving precision values ranging from 0.11 to 0.32 and recall values between 0.12 and 0.17. This translated to F1 scores between 0.12 and 0.16, with the highest performance observed in Jubba (IoU: 0.10) and NovaKakhovka (IoU: 0.11). The BCP method, while excelling in open areas, showed limited effectiveness in urban environments with precision below 0.005 and recall under 0.06 across all sites.
Table 4: Accuracy assessment of flood detection methods across three flood events for open areas.
@@
OTSU (VV) 0.03 0.45 0.05 0.03 0.09 0.52 0.15 0.08 0.04 0.63 0.07 0.04
OTSU (VH) 0.04 0.51 0.07 0.04 0.10 0.61 0.17 0.09 0.04 0.59 0.07 0.04
Siamese-CNN 0.38 0.53 0.44 0.28 0.42 0.32 0.36 0.22 0.22 0.13 0.16 0.09
BCP 0.78 0.84 0.81 0.68 0.63 0.80 0.70 0.54 0.68 0.36 0.47 0.31
Table 5: Accuracy assessment of flood detection methods across three flood events for urban areas.
@@
OTSU (VV) 0.00 0.23 0.00 0.00 0.00 0.09 0.00 0.00 0.00 0.02 0.00 0.00
OTSU (VH) 0.00 0.22 0.00 0.00 0.00 0.09 0.00 0.00 0.00 0.02 0.00 0.00
Siamese-CNN 0.11 0.17 0.12 0.08 0.32 0.12 0.14 0.10 0.26 0.17 0.16 0.11
BCP 0.01 0.06 0.01 0.01 0.00 0.03 0.00 0.00 0.00 0.02 0.00 0.00

### v3

@@ 4.2 Bayesian analysis for change point problems
$\begin{split}\frac{p_{i}}{1-p_{i}}&=\frac{P(U_{i}=1|\mathbf{X},U_{j},j\neq i)}{P(U_{i}=0|\mathbf{X},U_{j},j\neq i)}\\ &=\frac{\left[\int_{0}^{\gamma}p^{b}(1-p)^{n-b-1}dp\right]}{\left[\int_{0}^{\gamma}p^{b-1}(1-p)^{n-b}dp\right]}\cdot\frac{\left[\int_{0}^{\lambda}\frac{w^{b/2}}{(W_{1}+B_{1}w)^{(n-1)/2}}dw\right]}{\left[\int_{0}^{\lambda}\frac{w^{(b-1)/2}}{(W_{0}+B_{0}w)^{(n-1)/2}}dw\right]}\end{split}$ (1)
@@
We implemented BCP using the bcp R package [24]. We selected the default tuning parameters of $\gamma=\lambda=0.2$ as recommended by Barry and Hartigan [23, 24]. Posterior means and probabilities were calculated based on 500 MCMC iterations after discarding the first 50 as burn-in, a configuration that achieved convergence while maintaining computational efficiency.
@@
4.3 Post-processing
Speckle in SAR imagery results from the coherent summation of electromagnetic waves scattered from multiple targets within each resolution cell, creating a characteristic granular pattern that is an inherent property of coherent imaging systems. We applied the BCP to a Sentinel-1 GRD time series that inherently exhibits speckle patterns due to coherent imaging, which may lead to false detections of flooded areas. To mitigate speckle while preserving spatial details, we applied a moving window kernel [43, 44] to the results of the BCP analysis. The kernel that calculated the average of $w\times w$ pixel values centered on the target observation was applied to the BCP output. The window size $w$ is a parameter that can be adjusted based on the spatial resolution of the input imagery and the desired level of smoothing. After applying the filter, the smoothed values were thresholded using a parameter $t$. Pixels with values above this threshold were considered as potential flood areas, while those below were masked out.
To determine optimal parameters for our flood detection method, we conducted a systematic parameter sensitivity analysis focusing on the $w$ and $t$ as post-processing steps. For the window size parameter, we explored values ranging from 3 to 15 pixels with a step size of 2. The lower bound of 3 represents the minimum window size needed for effective speckle mitigation while the upper bound of 15 was chosen to avoid excessive smoothing that could eliminate small but significant flooded areas. The 20 m spatial resolution of our analysis data meant that a window size of 3 would cover an area of 60 m × 60 m on the ground, while a window size of 15 would cover 300 m × 300 m. For the threshold parameter $t$, we investigated values between 0.1 and 0.9 with a step size of 0.1. This comprehensive range covers the entire probability space from very permissive detection (0.1) to highly selective identification (0.9). Theoretically, changepoint probabilities above 0.5 indicate a more likely than not change, but in practice, lower thresholds may be effective for detecting subtle hydrological changes. The parameter selection directly influences accuracy assessment metrics through precision-recall trade-offs. Higher threshold values reduce both true and false positives, generally increasing precision but decreasing recall as actual flood areas may be missed. Larger window sizes reduce false detection through spatial smoothing, typically improving precision, though recall may decrease if small legitimate flood areas are also smoothed out. Understanding these relationships is crucial for interpreting the parameter optimization results and ensuring transparency in the accuracy assessment process. Thus, the parameter optimization process involved evaluating all 63 combinations of window sizes and thresholds (7 window sizes × 9 threshold values) across multiple flood events in diverse geographical settings to ensure the robustness and transferability of our approach.
@@ 5 Evaluation
5.0.1 Otsu’s Thresholding Method
As a conventional baseline, we implemented the Otsu’s threshold algorithm [45]. This non-parametric technique determines an optimal threshold by maximizing the between-class variance between flood and non-flood pixels in SAR imagery. We applied the Otsu’s thresholding approach to two different input configurations: VV polarization only and VH polarization only.
The Otsu’s threshold algorithm operates by computing the weighted sum of variances of the two classes:
@@ 5.0.1 Otsu’s Thresholding Method
where $\omega_{0}$ and $\omega_{1}$ represent the class probabilities, $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ denote the class variances, and $t$ is the threshold value. We followed standard practices in remote sensing literature for Otsu’s thresholding approach, including Lee filter application for speckle reduction, followed by histogram equalization preprocessing to enhance contrast in the SAR imagery before applying the threshold calculation.
@@ 5.0.2 Siamese CNN Method
The model was trained using binary cross-entropy loss function, with consideration for class imbalance where the positive class weight was calculated to be approximately 35.67. We used the Adam optimizer with an initial learning rate of 0.001 and gradient clipping ($clipnorm=1.0$), along with a ReduceLROnPlateau scheduler that reduced the learning rate by half after 5 epochs of no improvement. The model was trained for up to 200 epochs with early stopping ($patience=20$) based on validation loss. Each input was normalized using Z-score normalization and clipped to ±3 standard deviations to handle outliers. NaN values were replaced with zeros to ensure numerical stability.
@@ 5.0.3 Accuracy assessment
where $N_{TP}$ represents the number of true positives (correctly predicted flood water pixels), $N_{FP}$ represents the number of false positives (incorrectly predicted as flood water), and $N_{FN}$ represents the number of false negatives (flood water pixels not detected). The F1 score, which provides a balanced measure of the model’s performance, is calculated as:
@@
The IoU was aimed to quantify the spatial overlap between the predicted flood water regions and the ground truth annotations:
@@ 6.1 BCP Outputs
Figure 3: Randomly selected patches for: Jubba (Somalia) (a) Bing map that represent non-flood condition, (b) Referenced UrbanSARFloods data set, (c) BCP results used VV input only, (d) BCP results used VH input only, (e) BCP results used $V$-VH input; Weihui (China) (f) Bing map that represent non-flood condition, (g) Binary UrbanSARFloods data set, (h) BCP results used VV input only, (i) BCP results used VH input only, (j) BCP results used $V$-VH input; Nova Kakhovka (Ukraine) (k) Bing map that represent non-flood condition, (l) Binary UrbanSARFloods data set, (m) BCP results used VV input only, (n) BCP results used VH input only, (o) BCP results used VV-VH input.
@@ 6.2 Parameter Sensitivity Analysis
The original BCP results are the posterior probabilities of flood occurrence at each pixel. To delineate the flood extents, we post-processed the BCP results with different parameter settings for threshold and window size and created the binary (flood/non-flood) outputs. Considering the result using VV-VH input only, as this is thought to be the best performance for the BCP, Figure 4 shows the comparison of the binary outputs with different parameter settings for a selected patch of Jubba (Somalia) as an example. The larger $w$ tends to overestimate the flood extent, while the smaller $w$ may fail to mitigate speckle. The larger $t$ picks up only the flood extent with the higher confidence, while the smaller $t$ may include false positives.
Figure 4: Parameter sensitivity analysis for post-processing of BCP flood detection results using a randomly selected patch from Jubba, Somalia. The grid shows flood detection outputs for a representative 512×512 pixel patch across different combinations of filter window sizes (w = 3, 5, 7, 9, 11, 13, 15) and probability thresholds (t = 0.1 to 0.9). Blue areas indicate detected flood pixels, white areas represent non-flooded regions. The reference ground truth is shown in the leftmost column. F1 scores are displayed for each parameter combination to demonstrate performance variations.
@@
We identified 44 of 63 parameter combinations that showed no statistically significant difference from the optimal configuration ($p>0.05$). These combinations formed several key clusters: combinations with $t=0.2$ and window sizes 7, 11, and 13 (Mean F1 >> 0.59, see table 2); combinations with $t=0.3$ and window sizes 5-15 (Mean F1 >> 0.56); and some combinations with $t=0.1$ (Mean F1 >> 0.50). Performance degradation became statistically significant for threshold values $\geq$ 0.4, particularly when combined with extreme window sizes (Figure 5). The decline in performance showed a consistent pattern with increasing threshold values and window size.
Table 2: Top performing parameter combinations and their statistical significance against the result with parameters of t=0.2 and w=9. Selected results only with t=0.2.
@@ 6.3 Comparative Analysis Across Sites
With the use of the optimal parameters for BCP, Figure 6 illustrates the comparison of our method with the conventional and deep-learning approaches across randomly selected three example patches. Each column represents a different study area, with rows showing the referenced Ground Truth data, Otsu’s thresholding (VV, VH), Siamese CNN, and BCP methods. The Otsu’s thresholding approaches show poor performance across all sites, with significant false positives (overestimation) in Weihui and substantial false negatives (underestimation) in Jubba and Nova Kakhovka. Between the two Otsu’s thresholding variants, VV input appears to perform slightly better than VH input in terms of reducing scattered noise, though both struggle with accuracy. The Siamese CNN demonstrates considerably improved performance with reduced noise compared to the Otsu’s thresholding approaches, successfully eliminating much of the scattered false positives. However, it tends to underestimate the flood extent, particularly evident in Jubba and Nova Kakhovka where significant false negatives (red areas) are present. The BCP method shows the most promising results, achieving good flood detection with minimal noise while better capturing the full flood extent across all three study areas.
Figure 6: Comparative flood detection results across three example patches in Weihui (China), Jubba (Somalia), and Nova Kakhovka (Ukraine) using different methods. Rows show Ground Truth (UrbanSARFloods reference), Otsu’s thresholding (VV and VH polarizations), Siamese CNN, and the proposed BCP method.
@@ 6.4 Quantitative Performance Evaluation
The table 3 summarizes the accuracy measures of our approach compared to Otsu’s thresholding and Siamese CNN approaches. For this evaluation, we used the globally optimized parameters identified through sensitivity analysis ($t=0.2$, $w=9$) with VV-VH input. Our BCP approach achieved F1 scores ranging from 0.49 to 0.75, significantly outperforming Otsu’s thresholding and Siamese CNN approaches across all sites. Otsu’s thresholding approach struggled with false positives, resulting in low precision across all sites. While the Siamese CNN showed improved performance over Otsu’s thresholding approach, it still fell short of our method’s capabilities.
Table 3: Flood detection accuracy for Otsu’s thresholding, Siamese CNN, and Bayesian analysis of change point problems (BCP), evaluated across three flood events (Weihui, Jebba, and Nova Kakhovka), with the highest value in each column shown in bold.
@@
Otsu (VV) 0.01 0.35 0.03 0.01 0.07 0.54 0.12 0.07 0.02 0.58 0.04 0.02
Otsu (VH) 0.02 0.56 0.05 0.02 0.08 0.61 0.14 0.07 0.02 0.51 0.04 0.02
Siamese-CNN 0.72 0.23 0.35 0.21 0.90 0.26 0.40 0.25 0.72 0.23 0.35 0.21
BCP 0.71 0.78 0.75 0.59 0.60 0.78 0.68 0.52 0.52 0.46 0.49 0.36
@@ 6.5 Analysis of Performance Variations Between Urban and Open Areas
To better understand the performance characteristics across different land cover types, we conducted separate analyses for flood urban and flood open pixels. Tables 4 and 5 present the detailed accuracy metrics for each category and method. In open areas, all methods demonstrated measurable effectiveness in flood detection, though with varying degrees of success. The BCP method achieved the highest performance in Weihui with a precision of 0.71 and a recall of 0.78 (IoU: 0.59), followed by results in Jubba (precision: 0.60, recall: 0.78, IoU: 0.52). However, its performance declined significantly in Nova Kakhovka (precision: 0.59, recall: 0.46, IoU: 0.36). The Siamese CNN demonstrated a modest performance in open areas, particularly in Weihui where it achieved an IoU of 0.46, though its effectiveness varied considerably across sites. The Otsu’s thresholding approach showed a poor performance across sites, with recall values ranging from 0.33 to 0.61, though precision remained low (0.01-0.08).
The analysis of urban environments revealed more challenging conditions for flood detection across all methods. The conventional Otsu’s thresholding approach exhibited extremely low precision (0.00) despite achieving moderate recall values (0.63-0.88), resulting in poor overall performance with F1 scores (0.00). Similarly, the Siamese CNN demonstrated a poor performance in urban areas, with an F1 score and IoU of zero. The BCP method, while excelling in open areas, showed limited effectiveness in urban environments with precision below 0.05 and recall under 0.16 across all sites.
Table 4: Flood detection accuracy for Otsu’s thresholding, Siamese CNN, and Bayesian analysis of change point problems (BCP), evaluated across three flood events (Weihui, Jebba, and Nova Kakhovka) for open areas, with the highest value in each column shown in bold.
@@
Otsu (VV) 0.01 0.33 0.02 0.01 0.07 0.54 0.12 0.07 0.02 0.58 0.04 0.02
Otsu (VH) 0.02 0.56 0.04 0.02 0.08 0.61 0.14 0.07 0.02 0.51 0.04 0.02
Siamese-CNN 0.63 0.62 0.63 0.46 0.57 0.26 0.36 0.22 0.21 0.15 0.17 0.09
BCP 0.71 0.78 0.74 0.59 0.60 0.78 0.68 0.52 0.59 0.46 0.49 0.36
Table 5: Flood detection accuracy for Otsu’s thresholding, Siamese CNN, and Bayesian analysis of change point problems (BCP), evaluated across three flood events (Weihui, Jebba, and Nova Kakhovka) for urban areas, with the highest value in each column shown in bold.
@@
Otsu (VV) 0.00 0.70 0.00 0.00 0.00 0.88 0.00 0.00 0.00 0.69 0.00 0.00
Otsu (VH) 0.00 0.63 0.00 0.00 0.00 0.88 0.00 0.00 0.00 0.72 0.00 0.00
Siamese-CNN 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.02 0.00 0.00
BCP 0.05 0.13 0.07 0.03 0.00 0.03 0.00 0.00 0.03 0.16 0.05 0.03